Joe pays the house in 6 hours. Tom pays the house in 4 hours. Sue pays the house in 2 hours. If all of them worked together in how much time will the house be paid?
Answer by stanbon(75887)

Joe pays the house in 6 hours. Tom pays the house in 4 hours. Sue pays the house in 2 hours. If all of them worked together in how much time will the house be paid?
----------------
Joe rate = 1/6 house/hr
Tom rate = 1/4 house/hr
Sue rate = 1/2 house/hr
---------------------------
Together rate = 1/x house/hr
-----------------------------
Equation:
rate + rate + rate = together rate
1/6 + 1/4 + 1/2 = 1/x
2x + 3x + 6x = 12
11x = 12
x = 12/11 hr = 1 hr 5 min (time to do the house together)
=============================================================
Cheers,
Stan H.
-----------------